1. A computer company claims that its laptop batteries average more than 4.5 hours of use per charge with a standard deviation of 0.7 hours. A sample of 45 batteries last an average of 4.72 hours. At the .05 level of significance, is there evidence to support the claim that the batteries last more than 4.5 hours?
Answer #1

One sample z test :

Claim : Average battery life more than 4.5 hours

H0 : µ = 4.5   vs Ha : µ > 4.5  

Given : ar{X} = 4.72 , σ = 0.7 , n = 45

Population standard deviation σ is known therefore we use z statistic.

Test statistic:

Z =   rac{ar{X}-mu }{sigma /sqrt{n}}   = 4.72-4.5 0.7/V45

Z = 2.11

Critical value:     α = 0.05

As Ha contain > sign , this is right tail test,

So 1 - α = 0.95 therefore Z(1 -α ) = 1.645

Critical region: Reject H0, if Z ≥ 1.645 Or fail to reject H0 , if Z < 1.645

Decision: As Z is greater than 1.645 , we reject H0

Conclusion: There is significant evidence that average battery life more than 4.5 hours.